### v3

# Orthogonal projectors of LCD codes and their graph representations

Abstract
We establish one-to-one correspondences between (i) binary even LCD codes and certain simple graphs, and (ii) ternary LCD codes and certain two-graphs.
Keywords: LCD code, strongly regular graph.
Mathematical subject classification (2020): 94B05, 05C50.
1 Introduction
The interplay between codes and designs has provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is a symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied the relation between LCD codes and adjacency matrices of graphs. They gave two conditions under which adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such an adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. Furthermore, we prove an analogous results for the relation between ternary LCD codes and two-graphs. Haemers, Peeters, and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. This paper gives an explanation to the above observation.
The remaining part of this paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes, simple graphs, and two-graphs. We give our main results in Section 3. Section 4 establishes auxiliary results on the minimum weight of codes derived from strongly regular graphs, which are later used in Section 5. In that section, we apply our main results to analyze binary LCD codes obtained as row spans of adjacency matrices of strongly regular graphs. Finally, Section 6 presents concluding remarks and open problems.
2 Preliminaries
2.1 LCD codes
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. In particular, codes over $\mathbb{F}_{2}$ and $\mathbb{F}_{3}$ are said to be binary and ternary codes, respectively. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$, respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is a matrix whose rows form a basis of $C$.
The orthogonal projector $\Pi_{C}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is a unique linear map from $\mathbb{F}_{q}^{n}$ to $C$ satisfying
$v\Pi_{C}=\begin{cases}v&\textrm{if }v\in C,\\ 0&\textrm{otherwise}.\end{cases}$
An orthogonal projector is represented by a matrix $A$ satisfying $A^{T}=A$ and $A^{2}=A$. Conversely, any matrix satisfying $A^{T}=A$ and $A^{2}=A$ represents the orthogonal projector of some code. We refer the reader to [massey_linear_1992, bardet2019permutation] for more information on orthogonal projectors. In this paper, we often identify the orthogonal projector with its matrix representation in the standard basis.
The dual code $C^{\perp}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is defined as $C^{\perp}=\{x\in\mathbb{F}_{q}^{n}\mid(x,y)=0\text{ for all }y\in C\}$, where $(x,y)$ denotes the standard inner product. An $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is said to be an linear complementary dual code (LCD code for short) if $C\cap C^{\perp}=\{\mathbf{0}_{n}\}$. Massey [massey_linear_1992] introduced the concept of LCD codes, and gave the following characterization:
Theorem 2.1 (Proposition 1 [massey_linear_1992]).
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $G$ be a generator matrix of $C$. Then $C$ is an LCD code if and only if the $k\times k$ matrix $GG^{T}$ is nonsingular. Moreover, if $C$ is an LCD code, then $\Pi_{c}=G^{T}(GG^{T})^{-1}G$ is the orthogonal projector from $\mathbb{F}_{q}^{n}$ to $C$.
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. We will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva [bouyuklieva_optimal_2021] gave a characterization of both punctured codes and shortened codes of binary LCD codes:
Theorem 2.2 (Lemma 2 [bouyuklieva_optimal_2021]).
Let $C$ be a binary LCD $[n,k]$ code with $d(C),d(C^{\perp})\geq 2$. For all $1\leq i\leq n$, exactly one of $C^{i}$ and $C_{i}$ is an LCD code.
Theorem 2.3 (Proposition 2 [bouyuklieva_optimal_2021]).
Let $C$ be a binary even LCD $[n,k]$ code with $d(C),d(C^{\perp})\geq 2$. Then the punctured code of C on any coordinate is again an LCD code.
2.2 Strongly regular graphs
In this subsection, we review the definition and basic properties of strongly regular graphs. For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs]. A graph $\Gamma=(V,E)$ consists of a finite set of vertices $V$ together with a set of edges $E\subset\binom{V}{2}$. In this paper, we are concerned only with simple graphs, namely, graphs which are undirected, and without loops or multiple edges. We say that $x,y\in V$ are adjacent if $\{x,y\}\in E$, and denote by $x\sim y$. A vertex $y\in V$ is a neighbor of $x\in V$ if $x$ and $y$ are adjacent. The cardinality of the set of neighbors $N(x)=\{y\in V\mid x\sim y\}$ is called the valency of $x\in V$. A graph is $k$-regular if all its vertices have the same valency $k$. A simple regular graph is strongly regular with parameters $(v,k,\lambda,\mu)$ if it has $v$ vertices, valency $k$, and if any two adjacent vertices are together adjacent to $\lambda$ vertices, while any two nonadjacent vertices are together adjacent to $\mu$ vertices. A strongly regular graph with parameters $(v,k,\lambda,\mu)$ is denoted by $srg(v,k,\lambda,\mu)$. An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise. Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. The elementary counting argument gives the following theorem (see [cameron1991designs]).
Theorem 2.4.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$. Then $A^{2}=kI+\lambda A+\mu(J-I-A)$, where $I$ an $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively.
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of some strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained several computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. In particular, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
2.3 Two-graphs
In this subsection, we review the definition and basic properties of two-graphs. For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs]. A two-graph $(V,\Delta)$ consists of a set $V$ whose elements are called vertices, and a collection $\Delta$ of unordered triples of vertices such that each $4$-tuple of $V$ contains an even number of elements of $\Delta$ as subsets. Given a simple graph $\Gamma=(V,E)$, the set of triples of $V$ whose induced subgraph in $\Gamma$ contains an odd number of edges gives rise to a two-graph $(V,\Delta)$, and every two-graph can be represented in this way (see [cameron1991designs]).
Let us introduce another adjacency matrix of a simple graph called “$\mp 1$ adjacency matrix”: A $\mp 1$ adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=-1$ if $x\sim y$ and $A_{x,y}=1$ otherwise. It is known that simple graphs $\Gamma$ and $\Gamma^{\prime}$, with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$ respectively, represent the same two-graph if there is a diagonal $(1,-1)$-matrix $D$ such that $A^{\prime}=DAD$. This defines an equivalence class of graphs, called the switching class, and hence two-graphs can be identified with the switching class of simple graphs. The two-graphs $(V,\Delta)$ and $(V^{\prime},\Delta^{\prime})$ are isomorphic if there is a bijection $V\to V^{\prime}$ that induces a bijection $\Delta\to\Delta^{\prime}$. Let $\Gamma$ and $\Gamma^{\prime}$ be two simple graphs with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$, respectively. By definition, these simple graphs represent isomorphic two-graphs if there are a permutation matrix $P$ and a diagonal $(1,-1)$-matrix such that $A^{\prime}=P^{T}DADP$.
3 Main results
In this section, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. In what follows, we prove general results on orthogonal projectors of LCD codes over $\mathbb{F}_{q}$. Specific results on binary and ternary LCD codes will be stated in the following subsections.
Lemma 3.1.
Let $C$ be an LCD code over $\mathbb{F}_{q}$ with orthogonal projector $A$. Then $C$ equals the row span of $A$.
Proof.
Let $e_{i}$ be the unit vector whose components are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$. Then, $Ae_{i}=r_{i}\in C$ for all $1\leq i\leq n$. Hence the row span of $A$ is a subspace of $C$. On the other hand, for any $c\in C$ it holds that $Ac=c$. This tells us that any codeword $c$ can be expressed as a linear combination of the rows of $A$. Hence $C$ is a subspace of the row span of $A$. This completes the proof. ∎
Theorem 3.2.
Let $C$ and $C^{\prime}$ be LCD codes over $\mathbb{F}_{q}$ with orthogonal projectors $A$ and $A^{\prime}$, respectively. Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$.
Proof.
Let $G$ and $G^{\prime}$ be generator matrices of $C$ and $C^{\prime}$ respectively. If $C$ and $C^{\prime}$ are equivalent, then there are monomial matrices $M$ and $N$ such that $G^{\prime}=MGN$. Now the following computation shows that $A^{\prime}=N^{T}AN$.
$\begin{split}A^{\prime}&=(MGN)^{T}((MGN)(MGN)^{T})^{-1}(MGN)\\ &=N^{T}G^{T}M^{T}(MGG^{T}M^{T})^{-1}(MGN)\\ &=N^{T}G^{T}(GG^{T})^{-1}GN\\ &=N^{T}AN.\\ \end{split}$
Conversely, suppose that there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$. The row span of $A$ is preserved by both permuting coordinates and multiplying any coordinate by nonzero scalar. Thus multiplying $A$ by $N^{T}$ from left preserves the row span of $A$. Lemma 3.1 tells us that the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$, respectively. It follows that $C^{\prime}=CN$, and hence $C$ and $C^{\prime}$ are equivalent, as required. ∎
Remark 3.3.
An analogous statement about permutation equivalence was presented in [bardet2019permutation] in the context of public-key cryptography.
Lemma 3.4.
Let $C$ be an LCD code of length $n$ over $\mathbb{F}_{q}$ with orthogonal projector $A$, and let $r_{i}$ denote the $i$-th row of $A$. Then every diagonal entry of $A$ is zero if and only if $(r_{i},r_{i})=0$ for all $1\leq i\leq n$.
Proof.
This is straightforward for the $i$-th diagonal entry of $A^{2}$ equals $(r_{i},r_{i})$. ∎
3.1 Binary LCD codes and simple graphs
In this subsection, we restrict ourselves to binary codes. Our aim is to establish the equivalence between binary even LCD codes and certain simple graphs.
Lemma 3.5.
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some simple graph if and only if $C$ is an even code.
Proof.
If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero. By Lemma 3.4, this implies that any row of $A$ has even weight. Now, the result follows by Lemma 3.1.
Conversely, suppose $C$ is even. Since the orthogonal projector $A$ is a symmetric $(0,1)$-matrix, it remains to show that any diagonal entry of $A$ is 00. Since $C$ is even, any codeword of $C$ has even weight. In particular, every row of $A$ has even weight, and the result follows by Lemma 3.4. This completes the proof. ∎
Theorem 3.6.
The orthogonal projector $A$ of a binary even LCD code is an adjacency matrix of a simple graph satisfying $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Conversely, such an adjacency matrix is the orthogonal projector of a binary even LCD code. Furthermore, two binary even LCD codes are equivalent if and only if corresponding graphs are isomorphic.
Proof.
The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5. To prove the second part, recall that simple graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ respectively are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. Since any permutation matrix is a monomial matrix, Theorem 3.2 implies the result. ∎
Any adjacency matrix $A$ of a simple graph satisfying $A^{2}=A$ is the orthogonal projector of a binary LCD code, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. These graphs are characterized by the following conditions:
  • If two vertices $x$ and $y$ are adjacent, then the number of the common neighbors of $x$ and $y$ is odd.
  • If two vertices $x$ and $y$ are not adjacent, then the number of the common neighbors of $x$ and $y$ is even.
Among these graphs there are strongly regular graphs with parameters $(v,k,\lambda,\mu)$, where $k,\mu\equiv 0\pmod{2}$ and $\lambda\equiv 1\pmod{2}$.
Remark 3.7.
Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2 (see Section 2). To prove the theorem, they essentially considered the orthogonal projection of the unit vector $e_{i}$ and made the following observation: If the $i$-th component of the orthogonal projection of $e_{i}$ is $1$, then $C_{i}$ is LCD and $C^{i}$ is not LCD; otherwise, $C^{i}$ is LCD and $C_{i}$ is not LCD. The $i$-th component of the orthogonal projection of $e_{i}$ is equal to the $i$-th diagonal entry of the orthogonal projector $A$. Hence it follows that $(A)_{i,i}=0$ if and only if $C^{i}$ is LCD, and $A_{i,i}=1$ if only if $C_{i}$ is LCD.
3.2 Ternary LCD codes and two-graphs
In this subsection, we restrict ourselves to ternary codes. Our aim is to establish the equivalence between ternary LCD codes and certain two-graphs. Throughout this subsection, the term “adjacency matrix” means the $\mp 1$ adjacency matrix (see Section 2).
Lemma 3.8.
Let $C$ be a ternary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some two-graph if and only if every row of $A$ has weight divisible by three.
Proof.
If $A$ is an adjacency matrix of a two-graph, then any diagonal entry of $A$ is zero. The result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$.
Conversely, suppose that any row of $A$ has weight divisible by three. Since the orthogonal projector $A$ is a symmetric matrix whose entry is in $\{0,1,-1\}$, it remains to show that any diagonal entry of $A$ is 00. Again, the result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$. This completes the proof. ∎
In both binary and ternary cases, the conditions for orthogonal projectors to be adjacency matrices are characterized by weights of their row vectors. In binary case, Lemma 3.1 implies every codeword to have even weight, while an analogous statement is false in ternary case. In fact, if every codeword in $C$ has weight divisible by three, then $C$ is self-dual code and hence cannot be LCD. The condition “every row of the orthogonal projector has weight divisible by three” is not insightful since such a statement is seemingly not related to any property of the corresponding code. From this perspective, the following sufficient condition is more useful.
Theorem 3.9.
Let $C$ be a ternary LCD code of length $n$ with orthogonal projector $A$. If $A$ is an adjacency matrix of some two-graph, then the dimension of $C$ is divisible by three.
Proof.
In this proof, we regard $A$ as a matrix over $\mathbb{F}_{3}$. Since $A^{2}=A$, the eigenvalue of $A$ is either 00 or $1$. To see this, let $a$ be any eigenvalue of $A$ and $v$ be an eigenvector with respect to $a$. Then $0=(A^{2}-A)v=(a^{2}-a)v$, and hence $a\in\{0,1\}$, as required. Therefore, there is an invertible matrix $P$ over $\mathbb{F}_{3}$ such that $P^{-1}AP=\operatorname{diag}{(1,\dots,1,0,\dots,0)}$. Let $D$ denote this diagonal matrix. The code $C$ is exactly the $A$-invariant subspace, so its dimension equals the number of $1$s in $D$, which we denote by $k$. Now, the proof is concluded by the facts $\operatorname{tr}(A)=\operatorname{tr}(D)$ and $\operatorname{tr}(D)\equiv k\pmod{3}$, where in the latter we regard $\operatorname{tr}(D)$ as an integer. ∎
The converse of the above theorem is false: Let $C$ be a ternary code whose generator matrix is of the form $G=\begin{pmatrix}I_{k}&A\end{pmatrix},$ where $k\equiv 0\pmod{3}$ and $A$ is $k\times(n-k)$ matrix with $AA^{T}=O$. Since $GG^{T}=I_{k}$ is invertible, $C$ is LCD, and one can easily check that
$G^{T}(GG^{T})^{-1}G=\begin{pmatrix}I_{k}&A\\ A^{T}&O\end{pmatrix}.$
Because of the k $1$s in its diagonal, this matrix fails to be an adjacency matrix of any two-graph.
Theorem 3.10.
The orthogonal projector $A$ of a ternary LCD code is an adjacency matrix of a simple graph satisfying $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Conversely, such an adjacency matrix is the orthogonal projector of a ternary LCD code. Furthermore, two binary even LCD codes are equivalent if and only if corresponding graphs are isomorphic.
Proof.
The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5. To prove the second part, let $\Gamma$ and $\Gamma^{\prime}$ be two-graphs with adjacency matrices $A$ and $A^{\prime}$, and suppose their rows generate ternary LCD codes $C$ and $C^{\prime}$, respectively. If $\Gamma$ and $\Gamma^{\prime}$ are isomorphic, then there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $A^{\prime}=P^{T}DADP$. If we write $N=DP$, then $N$ is a monomial matrix and $A^{\prime}=N^{T}AN$. Hence $C$ and $C^{\prime}$ are equivalent by Theorem 3.2. Conversely, if $C$ and $C^{\prime}$ are equivalent, then there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$, again by Theorem 3.2. Recall that any monomial matrix can be represented by a diagonal matrix and a permutation matrix. Hence there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $N=DP$. Therefore, we have $A^{\prime}=P^{T}DADP$, as required. ∎
4 Bounds on minimum weights
In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in Section 5. Note that all assertions in this subsection hold not only for binary LCD codes but also for any binary codes.
Proposition 4.1.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A$. Then $d(C^{\perp})\geq 1+k/t$, where $t=\max\{\lambda,\mu\}$.
Proof.
If $d=d(C^{\perp})$, then there is a set of $d$ linearly dependent columns of $A$. Let $S$ be a submatrix of $A$ consists of such columns, and $n_{i}$ be the number of rows in $S$ with weight $i$. Since the sum of components of every row of $S$ equals to zero, the weight of every row of $S$ must be even. Counting in two ways the number of choices of $j$ 1s in each row of $S$ for $j=1,2$, we have
$\begin{split}\sum_{i}(2i)n_{2i}&=dk,\\ \sum_{i}(2i)(2i-1)n_{2i}&\leq d(d-1)t.\\ \end{split}$
It follows that $0\leq\sum_{i}2i(2i-2)\leq d((d-1)m-k)$, and hence $d\geq 1+k/m$. ∎
Proposition 4.2.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A+I$. Then $d(C^{\perp})\geq 1+(k+1)/(t+1)$, where $t=\max\{\lambda,\mu\}$.
Proof.
With the notation of that in the proof of Proposition 4.1, count in two ways the number of choices of $j$ 1s in each row of $S$ for $j=1,2$. Then we have
$\begin{split}\sum_{i}(2i)n_{2i}&=d(k+1),\\ \sum_{i}(2i)(2i-1)n_{2i}&\leq d(d-1)(t+1).\\ \end{split}$
It follows that $d\geq 1+(k+1)/(t+1)$. ∎
Corollary 4.3.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A$. If $k$ is odd, then $d(C^{\perp})$ is even.
Proof.
With the notation of that in the proof of Proposition 4.1, suppose that $d$ is odd. Counting in two ways the number of choices of 1s in each row of $S$, we have
$\sum_{i}(2i)n_{2i}=dk.$
However, this is absurd since the left hand side of the equation is even while the right hand side is odd. This completes the proof. ∎
5 Applications
Haemers, Peeters, and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. With the help of computers, they observed that
  1. (i)
    for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes;
  2. (ii)
    among such codes the code from the Paley graph had the largest minimum weight.
Based on these results, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]). In this section, we (i) prove that the former statement is true for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$, and (ii) verify the latter statement for $srg(41,20,9,10)$. Note that the result by Brouwer and Eijl [Brouwer1992-nc] shows that for strongly regular graphs $srg(25,12,5,6)$ and $srg(41,20,9,10)$, corresponding binary codes have dimension $12$ and $20$, respectively.
First, we verify that the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$. This can be done easily using the main results of this paper.
Corollary 5.1.
Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa.
Proof.
Let $A$ be an adjacency matrix of $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$. By Theorem 2.4 we have $A^{2}=\frac{q-1}{2}I+\frac{q-5}{4}A+\frac{q-1}{4}(J-I-A)$, where $I$ and $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively. Since we assume that $q\equiv 1\pmod{8}$, $A^{2}=A$ holds when we regard $A$ as a matrix over $\mathbb{F}_{2}$. This fact, together with Theorem 3.6, implies that the resulting code is an even LCD code. Hence the result follows. ∎
Next, we study the second part of the observation above. We remark that the classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights. On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet, making it difficult to prove the problem in a similar way.
Theorem 5.2.
The Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$.
Proof.
The result by Brouwer and Eijl [Brouwer1992-nc] shows that for $srg(41,20,9,10)$ any corresponding binary code has dimension $20$, respectively, and the Griesmer bound tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$. Recall that by the proof of Corollary 5.1, any $srg(41,20,9,10)$ gives an even LCD code, and hence the minimum weight must be at most $10$. A computer search verifies that the Paley graph gives a code which attains this bound. This concludes the proof. ∎
According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so we can simply use this fact to prove Theorem 5.2. However, the proof of this bound depends on a lengthy computer search, making it difficult to understand. From this perspective, the proof presented in this paper is more sophisticated. To the best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
Remark 5.3.
It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound. The author thinks that this problem is interesting.
According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism. We constructed binary even LCD codes from the $7152$ graphs, and verified that (i) there are mutually inequivalent; (ii) only the code that comes from the Paley graph has the largest minimum weight $10$; and (iii) all codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$. By Proposition 4.2, any code from $srg(41,20,9,10)$ must have minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.6, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
6 Concluding remarks
In this paper, we have established one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. Based on these results, we further studied the relation between LCD codes and graphs. Our motivation is to study whether a good graph gives a rise to a good code. We verified that for some strongly regular graphs the answer is positive, for (i) the automorphism groups of graphs equal those of the resulting codes, and (ii) the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all codes from $srg(41,20,9,10)$. It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound, and the author thinks that this problem is interesting.
7 Acknowledgement
The author would like to thank Stefka Bouyuklieva for her insightful questions and comments, which improved the manuscript.